what amount of ice that can be melted by 4000 calories of heat latent heat of fusion of ice =80 cal g-1

Grade:12th pass

1 Answers

jani
6 years ago
We know Q=m/l(latent heat of fusion of ice =80)Givem amount of heat(Q)=4000cal4000=m/80M=4000*80M=320000gM=320kg